The Taylor's plan to expand their 80 square foot garage by tripiling the dimensions. What will be the area of the new garage?
Applications of proportions
Thanks..Alex

Answer by drj(1380)

The Taylor's plan to expand their 80 square foot garage by tripling the dimensions. What will be the area of the new garage?

Step 1. Let w be the width and L be the length of the original dimensions.
Step 2. Original Area A1=wL=80

Step 3. Let 3w be tripled the width, and 3L be tripled the length since we are tripling the dimensions.

Step 4. New Area A2=3w*3L=9wL.

Step 5. A2=9wL=9*A1=9*80=720 square feet.

Step 6. The area of the new garage is 720 square feet.

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J